OpenStudy (anonymous):

the width of a rectangle is 12 units less than the length. the perimeter is 108 units. find the length

OpenStudy (anonymous):

hi can you help?

OpenStudy (anonymous):

2 widths and 2 lengths = 108 Therefore, 1 length + 1 width = 54 W = L-12 W = (54-12)/2 L = W+12

OpenStudy (anonymous):

is the answer 33

OpenStudy (anonymous):

Yup!! :-) Very good!!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Glad I could help! And thank you! :-)